Suppose that the demand and supply curves for ethanol in the United States are represented by the following equations:

QD = 1,600 − 320P
QS = −800 + 640P

where QD is the quantity demanded (in millions of gallons per month), QS is the quantity supplied, and P is the price (in dollars per gallon).

In the scenario above, if the market is in equilibrium, the price of ethanol is $(blank) per gallon and the quantity of ethanol sold is (blank) million gallons per month.

Now suppose that a severe drought in the Midwest raises the price of corn, the main ingredient used in ethanol production. As a result, the quantity of ethanol supplied decreases by 480 million gallons per month at every price. After the market has adjusted to the new conditions, the price of ethanol will be $(blank) per gallon and the quantity of ethanol sold will be (blank) million gallons per month.
Answer #1

At equilibrium, QD=QS

1600-320P= -800+640P

P=2.5

Q=800

Thus, the price of ethanol is $2.5 per gallon and the quantity of ethanol sold is 800 million gallons per month.

After a severe drought raises the price of corn implies supply decreases by 480 million per month hence supply curve changes to QS= -1280+640P

At equilibrium, 1600-320P = -1280+640P

P=3

Q=640